# EFPIX: A zero-trust encrypted flood protocol

### v1

@@ 3 Encoding
This concatenated data is used to generate a signature using the sender’s private key. The concatenated data is encrypted using the receiver’s public key which is also retrieved from the local (other alias) to (public key, my alias) map. The hash is computed on the encrypted blob and the signature. Finally, hash, nonce, encrypted blob and signature are concatenated to produce the final encoded message.
@@
Hash: It is used to determine whether the relaying nodes have already seen this message. It prevents infinite looping relays. A proof-of-work (PoW) requirement must be met to minimize spam and tampering similar to Bitcoin [6]. However, the nonce is excluded from the hash calculation. PoW difficulty is chosen to cap the global theoretical message rate, ensuring that even high-performance nodes cannot flood the network beyond a theoretical maximum. This cost is incurred only by the sender, making it a user-side performance trade-off.
@@ 4 Decoding and Relaying
Deduplication: The hash and nonce are extracted. Hash validity and the PoW is then verified. Every node keeps a record of hashes of previously relayed messages. If it has seen the hash, the message is discarded. If not, the hash is added to the record and the message is relayed to all neighbors.
@@ 6 Example Instantiation
Encoded message layout (580 bytes)
Version Hash Nonce Encrypted blob Signature
1 bytes 64 bytes 3 bytes 256 bytes 256 bytes

### v2

@@ 3 Encoding
This concatenated data is used to generate a signature using the sender’s private key. The concatenated data is encrypted using the receiver’s public key which is also retrieved from the local (other alias) to (public key, my alias) map. The hash is computed on the encrypted blob and the signature. Finally, nonce, encrypted blob and signature are concatenated to produce the final encoded message.
@@
Nonce: A proof-of-work (PoW) requirement must be met to minimize spam and tampering similar to Bitcoin [6]. However, the nonce is excluded from the hash calculation. PoW difficulty is chosen to cap the global theoretical message rate, ensuring that even high-performance nodes cannot flood the network beyond a theoretical maximum. This cost is incurred only by the sender, making it a user-side performance trade-off.
@@ 4 Decoding and Relaying
Deduplication: A hash is calculated on the packet excluding the nonce. It is used to determine whether the relaying nodes have already seen this message. It prevents infinite looping relays. The PoW is then verified. Every node keeps a record of hashes of previously relayed messages. If it has seen the hash, the message is discarded. If not, the hash is added to the record and the message is relayed to all neighbors.
@@ 6 Example Instantiation
Encoded message layout (516 bytes)
Version Nonce Encrypted blob Signature
1 bytes 3 bytes 256 bytes 256 bytes